### v2

# Uniform density in matroids, matrices and graphs

Abstract
We give new characterizations for the class of uniformly dense matroids and study applications of these characterizations to graphic and real representable matroids. We show that a matroid is uniformly dense if and only if its base polytope contains a point with constant coordinates. As a main application, we derive new spectral, structural and classification results for uniformly dense graphs. In particular, we show that connected regular uniformly dense graphs are $1$-tough and thus contain a (near-)perfect matching. As a second application, we show that uniform density can be recognized in polynomial time for real representable matroids, which includes graphs. Finally, we show that strictly uniformly dense real represented matroids can be represented by projection matrices with a constant diagonal and that they are parametrized by a subvariety of the Grassmannian.
1 Introduction
Matroids were introduced by Whitney and Nakasawa as an abstraction and generalization of independence in linear algebra (linear independence) and graph theory (acyclicity), and they have a rich combinatorial, geometric and representation theory; a standard reference is [24]. A matroid [^1] We note that this is one of many alternative ways, so-called cryptomorphisms, to define matroids. We refer to [24] for an overview of the most common definitions. $M=(E,\mathcal{B})$ is a nonempty collection $\mathcal{B}$ of subsets of a ground set $E$ that satisfy:
for all $\text{for all~{}}A,B\in\mathcal{B}\text{~{}and all $a\in A\backslash B$, thereexists $b\in B\backslash A$ such that $(A\backslash\{a\})\cup\{b\}\in\mathcal{B}$}.$ (1)
We will assume that every element of $E$ is contained in at least one element of $\mathcal{B}$; these are also called loopless matroids. Elements of $\mathcal{B}$ are called the bases of the matroid and an important consequence of property (1) is that every basis has the same cardinality, which is called the rank of the matroid. The rank can be extended to a function on subsets of the ground set as
$\operatorname{rank}(A):=\max\{|B\cap A|\,:\,\text{$B$ is a basis}\},\quad\text{for all $A\subseteq E$}.$
Following the terminology of [23], we call $\rho(A):=|A|/\operatorname{rank}(A)$ the density of a nonempty subset $A$, and call $\rho=\rho(M):=\rho(E)$ the density of the matroid. A matroid is called uniformly dense if $\rho(A)\leq\rho(E)$ for all nonempty subsets $A$.
Narayanan and Vartak [22] introduced uniformly dense matroids in the context of certain electrically-inspired partitions of graphs and matroids, and studied their basic properties in [23]. Catlin, Grossman, Hobbs and Lai [5] introduced the name “uniformly dense matroids” and showed that they are closely related to ‘fractional arboricity’ — this is the smallest $\alpha$ such that any collection of bases that covers every element of the ground set at least $t$ times, has size at least $\alpha\times t$ [25] — and ‘strength’ — this is the largest $\beta$ such that the rank of any subset of size $|E|-k$ is at least $\operatorname{rank}(E)-k/\beta$ [16, 7]. Catlin et al. showed that uniformly dense matroids are precisely those matroids for which strength and fractional aboricity coincide. Independently, uniformly dense matroids appeared in a conjecture on orderings of the bases of a matroid [17]; the conjecture says that a matroid $M$ is uniformly dense if and only if there exists an ordering of $E$ such that all $\operatorname{rank}(M)$ cyclically consecutive elements form a basis. Van den Heuvel and Thomassée [27] proved this conjecture for matroids whose size and rank are coprime.
The analysis of uniformly dense matroids in this article starts from the observation that the density inequalities $\rho(A)\leq\rho(E)$ can be interpreted geometrically as the hyperplane inequalities of a certain polyhedron, the base polytope $P(M)$ associated with the matroid $M$
$P(M)$ $:=\operatorname{conv}\left(\left\{e_{B}\,:\,\text{$B$ is a basis}\right\}\right)\subset\mathbb{R}^{|E|}$
=$=\left\{x\in\Delta_{M}\,:\,\sum_{e\in A}x_{e}\leq\operatorname{rank}(A)\text{~{}for all $\emptyset\neq A\subseteq E$}\right\},$ (2)
where $e_{B}$ is a 001111 indicator vector with entries $1$ for the elements in $B$, and with the hypersimplex $\Delta_{M}:=\left\{x\in\mathbb{R}^{|E|}_{\geq 0}\,:\,\sum_{e}x_{e}=\operatorname{rank}(M)\right\}$. Matroid polytopes were introduced by Edmonds [9] in the context of combinatorial optimization and independently by Gelfand, Goresky, MacPherson and Serganova [14] while studying cell decompositions of the Grassmannian. The halfspace description of $P(M)$ in (2) can be found in [10]. Our new characterization of uniformly dense matroids follows immediately from the definition of the matroid polytope.
Theorem 1.1.
The following are equivalent for a matroid $M$:
  1. (1)
    $P(M)$ contains the point $(\rho^{-1},\dots,\rho^{-1})$.
  2. (2)
    $\rho(A)\leq\rho(E)$ for all nonempty subsets $\emptyset\neq A\subseteq E$.
  3. (3)
    There exists a measure $\mu$ on $\mathcal{B}$ such that $\mu(\{B:B\ni e\})$ is equal for all $e\in E$.
A matroid that satisfies these conditions is called uniformly dense.
Proof.  $(1)\Leftrightarrow(2).$ The matroid polytope contains the point $(\rho^{-1},\dots,\rho^{-1})$ if and only if this point satisfies the hyperplane inequalities $\sum_{e\in A}\rho^{-1}\leq\operatorname{rank}(A)$, for all nonempty $A$. Equivalently,
$|A|\cdot\rho^{-1}\leq\operatorname{rank}(A)\iff\rho(A)=\frac{|A|}{\operatorname{rank}(A)}\leq\rho=\rho(E).$
$(1)\Leftrightarrow(3).$ The matroid polytope contains the point $(\rho^{-1},\dots,\rho^{-1})$ if and only if this point is a convex combination of the vertices of $P$. Equivalently, there must exist nonnegative coefficients $\mu(B)$ such that
$\sum_{B\in\mathcal{B}}\mu(B)e_{B}=(\rho^{-1},\dots,\rho^{-1})$
or, for each coordinate $e\in E$,
$\sum_{e\in B\in\mathcal{B}}\mu(B)=\rho^{-1}.$
These coefficients are the desired measure $\mu$ on $\mathcal{B}$. ∎
If the point $(\rho^{-1},\dots,\rho^{-1})$ in (1) lies in the relative interior of $P(M)$, then the matroid is called strictly uniformly dense. A measure that satisfies condition (3) in Theorem 1.1 will be called an $E$-uniform basis measure. The $E$-uniform basis measures associated with a uniformly dense matroid form a polytope: this is the intersection of the affine space
$\left\{\mu\in\mathbb{R}^{|\mathcal{B}|}\,:\,\sum_{B\ni e}\mu(B)=\rho^{-1},~{}\forall e\in E\right\}$
with the probability simplex. Several known results on uniformly dense matroids, for instance those in [5] related to fractional covering and packing, the conjecture of Kajitani et al. in [17] as well as our Theorem 4.6, can be interpreted as statements about certain points in this polytope.
As a first application of Theorem 1.1 we study graphic matroids, where $E$ are the edges of a graph $G$ and bases are spanning trees. The density of this matroid is denoted by $\rho(G)$ and a graph is said to be uniformly dense if the corresponding graphic matroid is uniformly dense. We study some implications of uniform density on the Laplacian spectrum of graphs and classify all connected uniformly dense graphic matroids with two cycles. Using $E$-uniform basis measures, we prove a connectivity result related to toughness: a graph $G$ is $t$-tough if removing any $k$ vertices yields at most $k/t$ disconnected components [2]. Any graph with maximum degree $\Delta(G)$ is $(1/\Delta(G))$-tough, and for uniformly dense graphs we find the following improvement: {restatable*}theoremtoughness A uniformly dense graph $G$ is $(\rho(G)/\Delta(G))$-tough. If, furthermore, $G$ is regular and connected, then it is $1$-tough.
This result also implies the existence of (near-)perfect matchings in regular uniformly dense graphs. The proof of Theorem 1.1 is based on techniques of Fiedler and is given in Section 3.3.
As a second application of Theorem 1.1 we consider real representable matroids, where $E$ is the set of columns of a real matrix and bases are maximal linearly independent sets of columns. For these matroids, we show in Theorem 4.6 that there exists an $E$-uniform measure of the form $\mu_{X}:B\mapsto\det(X_{B})^{2},$ given by the minors of some matrix $X$. As an important practical consequence of Theorem 1.1, we prove the following result on the computational complexity of recognizing uniform density in real representable matroids, which includes graphics matroids: {restatable*}theoremPolynomialtime Let $M$ be a real representable matroid. Given a representation of $M$, there is a polynomial-time algorithm that tests if $M$ is uniformly dense and, if it is not, outputs a density inequality that is violated.
A projection representation of a matroid is a projection matrix whose non-singular principal submatrices determine the bases of the matroid. Using results from the theory of determinantal point processes (see e.g. [18]) we show the following result in Section 4.2:
{restatable*}
theoremprojectionRepresentation Let $M$ be a real representable matroid. Then $M$ is strictly uniformly dense if and only if it has a projection representation with constant diagonal equal to $\rho(M)^{-1}$. As a final application, we show that strictly uniformly dense real representable matroids of size $n$ and rank $k$ can be parametrized by a subvariety $\mathcal{V}(n,k)$ of the Grassmannian, which parametrizes real representable matroids. This subvariety is further studied in follow-up work [8].
Paper outline. Section 2 recalls the definition of uniformly dense matroids, together with some examples. We describe basic properties and give some relevant known results.
Section 3 deals with uniformly dense graphic matroids. We introduce the required terminology and concepts from graph theory and give some examples. In Section 3.2 we study implications of uniform density on the structure of a graph and classify uniformly dense graphs with two cycles. Section 3.3 discusses connectivity of uniformly dense graphs with the toughness result Theorem 1.1 as the main result. Section 3.4 finally describes results on the normalized Laplacian spectrum of uniformly dense graphs.
Section 4 deals with real representable matroids. In Section 4.2 we consider strictly uniformly dense real representable matroids. We show that these matroids admit an $E$-uniform ‘determinantal’ measure in Theorem 4.6, that uniform density can be recognized efficiently in Theorem 1.1 and that they can be represented by a constant-diagonal projection matrix in Theorem 1.1. Section 4.3 introduces a variety that parametrizes strictly uniformly dense real matroids.
2 Uniformly dense matroids
2.1 Definition and examples
As introduced, a matroid $M=(E,\mathcal{B})$ consists of a nonempty set $\mathcal{B}$ that contains subsets of a finite ground set $E$, such that the base-exchange property (1) is satisfied. For convenience, we require that every element of the ground set is contained in at least subset in $\mathcal{B}$; these are also called loopless matroids (see later). Note that this implies $\operatorname{rank}(A)>0$ for every non-empty subset $A\subseteq E$ and excludes the empty matroid $(E,\emptyset)$. Elements $B\in\mathcal{B}$ are the bases of $M$ and a fundamental consequence of the base-exchange property is that every basis has the same cardinality. This is called the rank of $M$, and the rank of a set $A\subseteq E$ is defined as the size of the largest intersection of $A$ with any basis. A standard reference on matroid theory is [24]. We now define the ingredients for the definition of uniformly dense matroids.
Density: Narayanan and Vartak [22] defined the density $\rho(A):=|A|/\operatorname{rank}(A)$ of a nonempty set $A\subseteq E$; the density of a matroid is $\rho(M):=\rho(E)$. The possible densities of a matroid are $\rho\in\{|E|/1,\dots,|E|/|E|\}$ and Example 2.3 shows that every density is realized by some matroid.
Basis measure: A measure $\mu$ on the bases of $M$ is determined by a nonnegative real number $\mu(B)$ for each basis $B$. The measure is additive and we write $\mu(\mathcal{A})=\sum_{B\in\mathcal{A}}\mu(B)$ for the measure of a set $\mathcal{A}\subseteq\mathcal{B}$ of bases. A measure $\mu$ is normalized if $\mu(\mathcal{B})=1$, but this is not required in general. We note the following useful lemma on basis measures.
Lemma 2.1.
Every measure $\mu$ on $\mathcal{B}$ satisfies
$\sum_{e\in E}\mu(\{B:B\ni e\})=\operatorname{rank}(M)\mu(\mathcal{B}).$
Proof.
$\sum_{e\in B}\mu(\{B:B\ni e\})=\sum_{e\in E}\sum_{e\in B\in\mathcal{B}}\mu(B)=\sum_{B\in\mathcal{B}}\mu(B)|B|=\operatorname{rank}(M)\mu(\mathcal{B}).$
Base polytope: The base polytope $P(M)$ of a matroid is defined as the convex hull of indicator vectors of its bases $P(M)=\operatorname{conv}\left(\{e_{B}\,:\,B\text{~{}is a basis}\}\right)\subset\mathbb{R}^{|E|}.$ See Equation (2) or reference [10] for the description of this polytope in terms of hyperplane inequalities.
The following theorem from the introduction may be used as the definition of uniformly dense matroids. See 1.1
If in (1) the point lies in the relative interior of $P(M)$ or in (3) the measure has full support, then we say that $M$ is strictly uniformly dense. The second condition is a bit more subtle and is dealt with in Theorem 2.10 at the end of Section 2.2. We now give some examples of matroids and uniformly dense matroids.
Example 2.2 (Tadpole matroid).
The tadpole matroid $T_{4}$ is the matroid with ground set $[4]$ and bases 123, 134, 124, where strings denote subsets. The name “tadpole” is clarified in Example 3.3 and Figure 2. Since $\rho(234)=3/2>4/3=\rho(1234)$, the tadpole matroid is not uniformly dense. Theorem 2.5 will show that the fact that element $1$ is in every basis of $T_{4}$ is an obstruction for uniform density. The matroid polytope $P(T_{4})$ is the convex hull of the points
$\{(1,1,1,0),(1,1,0,1),(1,0,1,1)\}\subset\mathbb{R}^{4}.$
This is a regular triangle in the affine hyperplane $\{x\in\mathbb{R}^{4}:x_{1}=1\}$ and does not contain the point $(4/3,\dots,4/3)$.
Example 2.3 (Uniform matroid).
The uniform matroid $U_{n}^{k}$ of size $n$ and rank $k$ is the matroid with ground set $E=[n]:=\{1,\dots,n\}$ and bases =$\mathcal{B}=\{B\subseteq[n]:|B|=k\}=:{[n]\choose k}$ given by all $k$-element subsets of $[n]$. Since $\rho(U_{n}^{k})=n/k$, there exists a matroid with any rational density of at least 1. Let $\mu:B\mapsto 1$ be the uniform measure on the bases of $U_{n}^{k}$. Then $\mu(\{B:B\ni e\})={n-1\choose k-1}$ is independent of $e\in E$, which proves that uniform matroids are uniformly dense. The matroid polytope of a uniform matroid is equal to the associated hypersimplex $P(U_{n}^{k})=\Delta_{U_{n}^{k}}$ and contains the point $(k/n,\dots,k/n)$.
Example 2.4 (Self-dual matroids).
A matroid $M$ is self-dual if $B\in\mathcal{B}\Leftrightarrow E\backslash B\in\mathcal{B}.$ In particular, self-dual matroids have density $\rho(M)=2$. For any fixed basis $A$, the basis measure
$\mu_{A}:B\mapsto\{1/2\text{~{}if $B=A$~{}or~{}$B=E\backslash A$, and $0$ otherwise}\}$
is $E$-uniform, which proves that self-dual matroids are uniformly dense. Geiger, Hashimoto, Sturmfels and Vlad proved in [13] that the base polytope of self-dual matroids contains the point $(1/2,\dots,1/2)$. They called this property “stable” in reference to stability of point configurations.
2.2 Basic structure and operations
We start by giving a characterization of two forcing structures in uniformly dense matroids. An element $e\in E$ that is contained in every basis of a matroid is called a coloop and an element that is contained in no basis is called a loop; the following result shows that we do not lose much by restricting to loopless matroids when studying uniform density.
Theorem 2.5.
If a uniformly dense matroid contains a coloop (resp. loop), then all elements must be coloops (resp. loops).
Proof.  A coloop $\ell$ satisfies $\mu(\{B:B\ni\ell\})=\mu(\mathcal{B})$ for any basis measure $\mu$. If $M$ is uniformly dense with $E$-uniform basis measure $\mu$, this implies $\mu(\{B:B\ni e\})=\mu(\mathcal{B})$ for every $e\in E$, and by Lemma 2.1, that $\operatorname{rank}(M)=|E|$. This characterizes the uniform matroid $U_{n}^{n}=(E,E)$ in which every element is a coloop. A loop $\hat{\ell}$ satisfies $\mu(\{B:B\ni\hat{\ell}\})=0$ for any basis measure $\mu$. If $M$ is uniformly dense with $E$-uniform basis measure $\mu$, this implies $\mu(\{B:B\ni e\})=0$ for every $e\in E$ and by Lemma 2.1 that $\operatorname{rank}(M)=0$, which characterizes the empty matroid $U_{n}^{0}=(E,\emptyset)$ in which every element is a loop. ∎
Thus, a uniformly dense matroid of size $n$ is either $U_{n}^{n}$, empty or it has no loops or coloops.
Next, we show how uniform density behaves with respect to some common operations in matroid theory. Theorem 2.6 and 2.7 were proven by Narayanan and Vartak in [23], and Theorem 2.9 is new to our knowledge. These results are not used in later sections, but they are included to illustrate that uniformly dense matroids are a well-behaved class of matroids. The first operation is duality: the dual of a matroid $M=(E,\mathcal{B})$ is the matroid
$M^{\star}:=(E,\{E\backslash B:B\in\mathcal{B}\}).$
Theorem 2.6 ([23, Thm. 4]).
A matroid $M$ is uniformly dense if and only if its dual $M^{\star}$ is uniformly dense.
Proof.  We give a new proof based on our new characterization. Let $M$ be a matroid and $M^{\star}$ its dual. The bijection between $\mathcal{B}$ and $\mathcal{B}^{\star}$ via $B^{\star}=E\backslash B$ induces a bijection between measures $\mu$ on $\mathcal{B}$ and $\mu^{\star}$ on $\mathcal{B}^{\star}$. We find
$\sum_{e\in B\in\mathcal{B}}\mu(B)=\sum_{e\in B\in\mathcal{B}}\mu^{\star}(E\backslash B)=\sum_{e\notin B^{\star}\in\mathcal{B}^{\star}}\mu^{\star}(B^{\star})=\mu^{\star}(\mathcal{B}^{\star})-\sum_{e\in B^{\star}\in\mathcal{B}^{\star}}\mu^{\star}(B^{\star}),$
where the last step uses additivity of the measure over the partition =$\mathcal{B}^{\star}=\{B^{\star}:B^{\star}\ni e\}\sqcup\{B^{\star}:B^{\star}\not\ni e\}.$ Since $\mu^{\star}(\mathcal{B}^{\star})$ is independent of $e$, the measure $\mu$ is $E$-uniform if and only if $\mu^{\star}$ is $E$-uniform and thus $M$ is uniformly dense if and only $M^{\star}$ is. ∎
The union of two matroids $M_{1}=(E_{1},\mathcal{B}_{1})$ and $M_{2}=(E_{2},\mathcal{B}_{2})$ is the matroid
$M_{1}\lor M_{2}:=(E_{1}\cup E_{2},\max\{B_{1}\cup B_{2}:B_{1}\in\mathcal{B}_{1},B_{2}\in\mathcal{B}_{2}\}),$
where ‘$\max$’ is taken with respect to the subset partial order.
Theorem 2.7 ([23, Cor. 1]).
The union of two uniformly dense matroids $M_{1}$ and $M_{2}$ on the same ground set $E$ is a uniformly dense matroid $M_{1}\lor M_{2}$ on $E$.
The following example shows that the converse of Theorem 2.7 does not hold: the matroid $M=([3],\{12,13\})$ is not uniformly dense but the union $M\lor M=U_{3}^{3}$ is uniformly dense.
The intersection of two matroids $M_{1}$ and $M_{2}$ is the matroid
$M_{1}\land M_{2}:=(M_{1}^{\star}\lor M_{2}^{\star})^{\star}.$
It follows immediately from Theorem 2.6 and 2.7 that intersections preserve uniform density.
Corollary 2.8.
The intersection of two uniformly dense matroids $M_{1}$ and $M_{2}$ on the same ground set is a uniformly dense matroid $M_{1}\land M_{2}$ on $E$.
The direct sum of two matroids $M_{1}=(E_{1},\mathcal{B}_{1})$ and $M_{2}=(E_{2},\mathcal{B}_{2})$ with disjoint ground sets is the matroid
$M_{1}\oplus M_{2}:=(E_{1}\cup E_{2},\{B_{1}\cup B_{2}:B_{1}\in\mathcal{B}_{1},B_{2}\in\mathcal{B}_{2}\}).$
In other words, the direct sum is the union of matroids on disjoint ground sets.
Theorem 2.9.
The direct sum $M_{1}\oplus M_{2}$ of two matroids is uniformly dense if and only if $M_{1}$ and $M_{2}$ are uniformly dense with equal density $\rho(M_{1})=\rho(M_{2})=\rho(M_{1}\oplus M_{2})$.
Proof.  From the definition of direct sum, it follows that the base polytope of a direct sum matroid is the cartesian product of the base polytopes of the summands, that is $P(M_{1}\oplus M_{2})=\{(x,y)\in\mathbb{R}^{|E_{1}|+|E_{2}|}:x\in P(M_{1}),y\in P(M_{2})\}$. Hence, $P(M_{1}\oplus M_{2})$ contains the constant vector $\rho(M_{1}\oplus M_{2})\cdot\mathbbm{1}$ if and only if the polytopes $P(M_{1})$ and $P(M_{2})$ contain the constant vectors $\rho(M_{1}\oplus M_{2})\cdot\mathbbm{1}$ of appropriate lengths. This is possible if and only if $\rho(M_{1}\oplus M_{2})$ equals $\rho(M_{1})$ and $\rho(M_{2})$ since the base polytopes $P(M_{1})$ and $P(M_{2})$ lie in the hyperplane of vectors which sum to the rank of the matroid, so if these polytopes contain a constant vector then it must have values equal to the density of the matroid. This completes the proof. ∎
We now come back to the rank-based characterization of strict uniform density. A matroid is connected if it cannot be written as the direct sum of two other matroids. Every matroid $M$ has a partition $E=E_{1}\sqcup\dots\sqcup E_{\ell}$ of its ground set such that $M=\bigoplus_{i=1}^{\ell}M_{i}$ where each summand $M_{i}$ is a connected matroid on $E_{i}$; these are called the connected components of $M$.
Theorem 2.10.
The following are equivalent for a matroid $M$:
  • (1)
    $P(M)$ contains the point $(\rho^{-1},\dots,\rho^{-1})$ in its relative interior.
  • (2)
    $\rho(A)\leq\rho(E)$ for all nonempty subsets $\emptyset\neq A\subseteq E$, with strict inequality whenever $A$ is not the union of ground sets of connected components of $M$.
  • (3)
    There exists a positive measure $\mu$ on $\mathcal{B}$ such that $\mu(\{B:B\ni e\})$ is equal for all $e\in E$.
A matroid that satisfies these conditions if called strictly uniformly dense.
Proof.  The equivalence between (1) and (3) follows as in the proof of Theorem 1.1. To prove equivalence with (2), we use the fact, shown in [10, Prop. 2.4], that $P(M)$ has codimension equal to the number of connected components of $M$. Let $M=\bigoplus_{i=1}^{\ell}M_{i}$ be a decomposition into connected components. From the definition of direct sum it follows that $\sum_{e\in E(M_{i})}x_{e}=\operatorname{rank}(E(M_{i})$ for all $x\in P(M)$ and for every connected component $M_{i}$. These are $\ell$ linearly independent equations for points in the base polytope $P(M)$. Since the codimension of $P(M)$ is $\ell$, any further independent face inequality $\sum_{e\in A}x_{e}\leq\operatorname{rank}(A)$ must determine a proper face of $P(M)$. Thus the constant point $(\rho^{-1},\dots,\rho^{-1})$ lies in the relative interior if and only if these proper face inequalities are strict, which translates to $\rho(A)<\rho(E)$ following the proof of Theorem 1.1, whenever $A$ is not a union of some of the $E(M_{i})$’s. This completes the proof. ∎
3 Graphic matroids
We now consider matroids which are associated with graphs and study some implications of uniform density for the structure of these graphs. The main results are Theorem 3.10, which characterizes bicyclic connected uniformly dense graphs, Theorem 1.1, which gives a strong connectivity result for uniformly dense graphs and Theorem 3.14, which gives a spectral characterization of uniformly dense graphs.
3.1 Definition and examples
A graph $G=(V,E)$ consists of a finite set $V$ and a set [^2] If $E$ is allowed to be a multiset, with possible repeated edges, then the associated graph is called a multigraph. While this definition is most natural in the context of matroids, we only consider the special case of multigraphs without edge repeats. These are also called simple graphs or graphs. $E\subseteq\{\{u,v\}:u,v\in V\}$ of $2$-element subsets of $V$. Elements of $V$ are called vertices and elements of $E$ are called edges. For two distinct edges $e$ and $f$, we write $e\sim f$ if $e\cap f\neq\emptyset$. A path in $G$ is an ordered set of edges $e_{1}\sim e_{2}\sim\dots\sim e_{\ell}\subseteq E$ such that every vertex appears at most once in the intersection $(e_{k}\cap e_{k+1})_{1\leq k<\ell}$ between consecutive edges, and a cycle in $G$ is a path where $e_{1}\sim e_{\ell}$. A spanning tree in $G$ is a maximal subset of edges $T\subseteq E$ which contains no cycle [^3] We note that a spanning tree $T$ need not induce a connected graph $(V,T)$.. The collection of spanning trees of a graph determines a matroid
$M(G)=(E,\{T\subseteq E\,:\,T\text{~{}is a spanning tree in $G$}\}).$
$M(G)$ is called the cycle matroid of $G$, and any matroid which can be realized as the cycle matroid of a (multi-)graph is called a graphic matroid. In general, a single matroid can be realized by different graphs (as we shall see for instance in Example 3.4).
The rank of a graph is $\operatorname{rank}(G)=\operatorname{rank}(M(G))$, and the rank of a subset of edges is defined as the rank of the corresponding subset in $M(G)$. Graph theoretically, the rank of a graph is related to the number of components and the number of cycles in the graph. The number of components $c(A)$ of a subset $A\subseteq E$ is defined as
$c(A):=|V|-\operatorname{rank}(A),$
and we write $c(G)$ for the number of components of a graph [^4] The number of components $c(G)$ of a graph does not correspond to the number of connected components of the matroid $M(G)$. Instead, the number of connected components of $M(G)$ equals the number of biconnected components of $G$ (see [24, Prop. 4.1.7]); a biconnected component is a maximal subgraph without cut vertices.. This terminology reflects that $c(A)$ is the number of equivalence classes in the graph $(V,A)$ with respect to the equivalence relation $e\equiv f$ if and only if there exists a path with ends $e$ and $f$. A connected graph is a graph $G$ with $c(G)=1$. The number of cycles in a subset $A\subseteq E$ is defined as
$\beta(A):=|A|-\operatorname{rank}(A),$
and equals the number of linearly independent cycles in the set $A$; we write $\beta=\beta(G)$ for the number of cycles in a graph. The density $\rho(G):=\rho(M(G))$ of a graph can thus be expressed as
$\rho(G)=\frac{|E|}{\operatorname{rank}(G)}=\frac{|E|}{|V|-c(G)}=\frac{|E|}{|E|-\beta(G)}.$ (3)
A uniformly dense graph is a graph $G$ whose cycle matroid $M(G)$ is uniformly dense. Since the cycle matroid of a graph is the direct sum of the cycle matroids of the connected components of the graph, it follows from Theorem 2.9 that a graph is uniformly dense if and only if each connected component is uniformly dense with equal density. In the rest of this section, we study properties of uniformly dense graphs.
Remark 3.1.
The name “density” for matroids corresponds to what is classically known as density for graphs. A graph $G$ is called dense if the number of edges is close to the maximal possible number of edges, and it is called sparse otherwise. Hence, the complete graph is the most dense graph, and the further $G$ is from the complete graph, the sparser it is. Also, the denser $G$ is, the larger $|E|$ is compared to $|V|$, and therefore the larger $\rho(G)$ is, following (3). This gives the following interpretation of uniform density in graphs: a graph $G$ is uniformly dense if, for all nonempty $A\subseteq E$, the graph $(V,A)$ is at most as dense as $G$.
Example 3.2.
Figure 1 shows a graph $G$ with $|V|=4$ vertices and $|E|=5$ edges. The graph is connected, that is, $c(G)=1$, and thus it has $\operatorname{rank}(G)=|V|-1=3$. On the right, all spanning trees are shown — these are the sets of $3=\operatorname{rank}(G)$ edges that cover all vertices. The cycle matroid $M=M(G)$ corresponding to $G$ has ground set $E(M)=[5]$, given by the labeling of the edges of $G$, and $8$ bases $\mathcal{B}(M)=\{123,234,134,124,145,245,135,235\}$ given by the spanning trees of $G$. The non-bases of $M$ are $\{125,345\}$ and correspond to the $3$-cycles in the graph.
Figure 1: A graph $G$ and its spanning trees. The cycle matroid of $G$ has ground set $E=[5]$ and bases given by the three-element subsets indicated under the spanning trees.
Example 3.3 (Tadpole graph).
The tadpole graph (also called paw graph) shown in Figure 2 looks like a tadpole and has the tadpole matroid $T_{4}$ as cycle matroid. As shown in Example 2.2, the tadpole graph is not uniformly dense.
Example 3.4 (Tree graph).
A tree graph [^5] A tree with multiple connected components is also called a forest. is a graph without cycles and the cycle matroid of a tree graph with $n$ edges is the uniform matroid $U_{n}^{n}$. This illustrates how one matroid ($U_{n}^{n}$) can correspond to multiple graphs (all tree graphs on $n$ edges). Since the uniform matroid is uniformly dense, tree graphs are uniformly dense. A cut edge in a graph is an edge $e$ such that $c(E\backslash e)=c(E)+1$ and corresponds to a coloop in the cycle matroid. Following Theorem 2.5, a uniformly dense graph is either a tree graph or it contains no cut edges. A loop in a graph is an edge of the form $\{v,v\}$ and corresponds to a loop in the cycle matroid. Loops are excluded in our definition of graphs, but this is without much loss of generality when studying uniformly dense graphs due to Theorem 2.5.
Example 3.5 (Edge-transitive graphs).
An automorphism of a graph $G=(V,E)$ is a permutation $\pi$ of $V$ such that
$\{\pi(u),\pi(v)\}\in E\iff\{u,v\}\in E.$
The automorphisms of a graph form a group $\operatorname{Aut}(G)$ and a graph is edge transitive if $\operatorname{Aut}(G)$ acts transitively on $E$ via $\pi\{u,v\}=\{\pi(u),\pi(v)\}$. In other words, $G$ is edge transitive if for all $e,f\in E$ there exists an automorphism $\pi\in\operatorname{Aut}(G)$ such that $\pi(e)=f$. Narayanan and Vartak showed in [23] that edge-transitive graphs are uniformly dense. A proof follows from the observation that the uniform measure on the bases of an edge-transitive graph is $E$-uniform. Figure 2 shows some common examples of edge-transitive graphs: the cycle graph, complete graph and hypercube graph.
Figure 2: From left to right: (Example 3.3) The tadpole graph has density $\rho=4/3$ and is not uniformly dense. (Example 3.4) The tree graph has density $\rho=1$ and is uniformly dense. (Example 3.5) The $4$-cycle graph has density $\rho=4/3$ and is uniformly dense, the complete graph on $4$ vertices has density $\rho=2$ and is uniformly dense and the hypercube graph of dimension $3$ has density $\rho=12/7$ and is uniformly dense. The latter three graphs are edge-transitive.
3.2 Structure
To illustrate the implications of uniform density in graphs, we consider a number of structural properties. The degree $\deg(v):=|\{e\in E:e\ni v\}|$ of a vertex $v\in V$ is equal to the number of edges that contain $v$. We write $\delta(G)$ for the smallest degree of any vertex, and $\Delta(G)$ for the largest degree. A graph is regular if all vertices have the same degree.
Proposition 3.6.
If $G$ is a uniformly dense graph, then $\delta(G)\geq\rho(G)$.
Proof.  Let $v\in V$ be a minimal degree vertex, and let $A_{v}:=\{e\in E:e\ni v\}$ be the set of edges that contain $v$. Note that $|A_{v}|=\delta(G)$. Since removing $A_{v}$ creates at least one new component, i.e., the separated vertex $v$, we find that
$\operatorname{rank}(E\setminus A_{v})=|V|-c(E\setminus A_{v})\leq|V|-c(E)-1\Rightarrow\operatorname{rank}(E)-\operatorname{rank}(E\setminus A_{v})\geq 1.$
From uniform density of $G$, it then follows that
$\rho(E\setminus A_{v})=\frac{|E|-\delta(G)}{\operatorname{rank}(E\setminus A_{v})}\leq\frac{|E|}{\operatorname{rank}(E)}\iff\delta(G)\geq\rho(G)[\operatorname{rank}(E)-\operatorname{rank}(E\setminus A_{v})]\Rightarrow\delta(G)\geq\rho(G)$
which completes the proof ∎
A clique in a graph $G$ is a set of vertices $U\subseteq V$ such that $\{u,v\}\in E$ for every pair of distinct vertices $u,v\in U$. The clique number $\operatorname{cl}(G)$ is the largest number of vertices in a clique in $G$.
Proposition 3.7.
If $G$ is a uniformly dense graph, then $\operatorname{cl}(G)\leq 2\rho(G)$.
Proof.  Let $G$ be a uniformly dense graph and $A$ the edges in a maximal clique of size $k=\operatorname{cl}(G)$. Then
$\rho(E)\geq\rho(A)=\frac{|A|}{|V|-c(A)}=\frac{\frac{k(k-1)}{2}}{k-1}=\frac{k}{2}.$
The girth $\operatorname{gir}(G)$ of a graph $G$ is the size of the smallest cycle in $G$.
Proposition 3.8.
If $G$ is a uniformly dense graph with a cycle, then $\operatorname{gir}(G)\geq\rho(G)/(\rho(G)-1)$.
Proof.  Let $G$ be a uniformly dense graph with a cycle, and let $A$ be the set of edges in a smallest cycle of size $k=\operatorname{gir}(G)$. Then,
$\rho(E)\geq\rho(A)=\frac{|A|}{|V|-c(A)}=\frac{k}{k-1}.$
This simplifies to $k\geq\rho(G)/(\rho(G)-1)$ and completes the proof. ∎
The definition of girth extends to matroids. A circuit is a subset $C\subseteq E$ such that $\operatorname{rank}(C)=|C|-1$ and $\operatorname{rank}(S)=|S|$ for every proper subset $S\subset C$; the girth $\operatorname{gir}(M)$ of a matroid is the size of the smallest circuit. The proof of Proposition 3.8 directly extends to general matroids:
Proposition 3.9.
If $M$ is uniformly dense with a circuit, then $\operatorname{gir}(M)\geq\rho(G)/(\rho(G)-1)$.
When graphs have a small number of cycles, there is a strong interplay between the lengths of these cycles and the density of the graph. As an application, we get the following characterization for connected uniformly dense graphs with two cycles.
Theorem 3.10.
Let $G$ be a connected graph with $\beta(G)=2$. Then $G$ is uniformly dense if and only if $G$ can be constructed from three paths identified at the ends, with lengths $L_{1},L_{2}$ and $L_{3}$ that satisfy $(L_{3}-L_{2})\leq L_{1}\leq L_{2}\leq L_{3}$.
In the case where one of the paths has length zero, the graph in Theorem 3.10 consists of two cycles of equal length identified in a single vertex (see Example 3.11). In the proof below, the forward direction was suggested by Harry Richman and simplifies an earlier version.
Proof.  Note that since $G$ is connected, $\beta(G)=2$ implies that $|E|=|V|+1$.
(Proof of forward direction.) Let $G$ be a graph constructed from three paths identified at the ends. Let $P_{1},P_{2},P_{3}$ denote the edge sets of the paths, and let $L_{1},L_{2},L_{3}$ be their respective lengths. We show that the dual matroid $M(G)^{\star}$ is uniformly dense if and only if the inequalities of the theorem hold. Since $G$ is a planar graph, the dual matroid $M(G)^{\star}$ is again a graphic matroid, with corresponding graph given by the planar dual $G^{\star}$ of $G$ (see [24, Thm. 2.3.4]); Figure 3 shows an example. The graph $G^{\star}$ is a multigraph on $3$ vertices, say $\{a,b,c\}$ with $L_{1}$ edges $E_{ab}$ between $a$ and $b$, $L_{2}$ edges $E_{bc}$ between $b$ and $c$ and $L_{3}$ edges $E_{ac}$ between $a$ and $c$; see the figure below for illustration. $M(G)^{\star}$ has rank two and its bases are given by pairs of edges that do not lie between the same vertices.
Figure 3: A planar graph $G$ consisting of paths of length $L_{1}=3,L_{2}=3,L_{3}=5$. The three vertices of the dual graph $G^{\star}$ correspond to the two bounded and one unbounded face of $G$.
We now show that the path-length inequalities imply the density inequalities for all non-empty subsets $A\subset E=E_{ab}\sqcup E_{bc}\sqcup E_{ac}$. The subsets $A$ of rank $2$ are obtained by taking a basis and adding any further elements; in this case, $\rho(A)=|A|/2\leq|E|/2=\rho(E)$. All remaining nonempty subsets $A$ have rank $1$ and are obtained by taking a non-empty subset of $E_{ab},E_{bc}$ or $E_{bc}$. Since the density satisfies $\rho(A)=|A|$, the extreme cases for which the density inequalities need to be checked correspond to when $A$ is equal to one of the three sets of parallel edges, in these cases $\rho(E_{ab})=L_{1}$, $\rho(E_{bc})=L_{2}$ and $\rho(E_{ac})=L_{3}$. We check the density inequalities:
$\rho(E_{ab})$ $\leq\rho(E)\iff 2L_{1}\leq L_{1}+L_{2}+L_{3}\iff L_{1}\leq L_{2}+L_{3}\Longleftarrow L_{1}\leq L_{2}$
$\rho(E_{bc})$ $\leq\rho(E)\iff 2L_{2}\leq L_{1}+L_{2}+L_{3}\iff L_{2}\leq L_{1}+L_{3}\Longleftarrow L_{2}\leq L_{3}$
$\rho(E_{ac})$ $\leq\rho(E)\iff 2L_{3}\leq L_{1}+L_{2}+L_{3}\iff L_{3}\leq L_{1}+L_{2}\Longleftarrow L_{1}\geq(L_{3}-L_{2}).$
Thus, uniform density of $M(G)^{\star}$ follows from the assumptions on the path lengths. This completes the forward direction.
(Proof of converse direction.) Let $G$ be a connected uniformly dense graph with $|E|=|V|+1$ edges. We will show that $G$ can be constructed from three paths identified at the ends. The conditions on the lengths then follow from the proof of the forward direction above.
Let $T$ be any spanning tree of $G$. This tree is connected since $G$ is connected and, by definition of rank, it has $|T|=\operatorname{rank}(G)=|E|-2$ edges. The graph $G$ is thus constructed from a connected tree to which two edges are added. By Proposition 3.6 and $\rho(G)>1$, the minimal degree of $G$ is at least two and thus $G$ has no leaves. This means that the tree $T$ can have at most four leaves and that these must be contained in the added edges when going from $T$ to $G$.
Figure 4: $T_{1},T_{2},T_{3},T_{4}$ are the four possible trees with at most four leaves — they are drawn here without degree-two vertices. Next to each tree we show the possible ways to add two edges such that no degree-one vertices remain and cut edges are indicated in red.
Figure 4 shows the possible trees with at most four leaves — only vertices with degree different from two are drawn — and the possible ways of adding two edges to these trees to obtain a graph without leaves. Since $G$ is uniformly dense and not a tree, the graphs with cut edges (coloops) are excluded by Theorem 2.5. Figure 4 thus enumerates all possibilities for $G$ and all these possibilities can be constructed from three paths identified at the ends. This completes the proof. ∎
Example 3.11.
Let $G,H,K$ be the graphs in Figure 5. These graphs are constructed by identifying the ends of three paths of lengths $(0,4,4)$ for $G$, lengths $(3,3,6)$ for $H$ and lengths $(2,3,6)$ for $K$. Applying Theorem 3.10, we find that $G$ is uniformly dense (since $4-4\leq 0\leq 4\leq 4$) and $H$ is uniformly dense (since $6-3\leq 3\leq 3\leq 6$) but $K$ is not uniformly dense (since $6-3\not\leq 2\leq 3\leq 6$).
Figure 5: The three graphs $G,H$ and $K$ from Example 3.11. $G$ and $H$ are uniformly dense and $K$ is not uniformly dense.
3.3 Connectivity
Uniform density has strong implications for the connectivity of a graph. Recall that $c(A)$ is the number of connected components of the graph $(V,A)$ and thus that $c(E\backslash A)-c(E)$ is the number of new components created when removing a set of edges $A$ from the graph. The uniform density condition of a graph can be written in terms of the number of connected components as follows:
Proposition 3.12.
A graph $G$ is uniformly dense if and only if
$c(E\backslash A)-c(E)\leq\frac{|A|}{\rho(G)}\text{~{}for all $A\subset E$.}$ (4)
Proof.  Let $G$ be a graph and $A\subset E$ a proper subset of edges. Then,
$\rho(E\backslash A)\leq\rho(E)$ $\iff\frac{|E\backslash A|}{\operatorname{rank}(E\backslash A)}\leq\frac{|E|}{\operatorname{rank}(E)}$
$\iff\frac{|E|-|A|}{|V|-c(E\backslash A)}\leq\frac{|E|}{|V|-c(E)}$
$\iff c(E\backslash A)-c(E)\leq|A|\frac{|V|-c(E)}{|E|}$
$\iff c(E\backslash A)-c(E)\leq\frac{|A|}{\rho(G)}.$
Since $E\backslash A$ can be any nonempty set of edges, this completes the proof. ∎
In other words, removing $k$ edges from a uniformly dense graph increases the number of components by at most $k/\rho(G)$. This notion of connectivity is also called “strength” and its reciprocal “vulnerability” and was studied by Gusfield in [16] for graphs, by Cunningham in [7] for matroids. Strength was related to uniform density by Catlin et al. in [5].
[…]
$n_{U}\leq c(E\backslash A_{U})-c(E)\leq\frac{|A_{U}|}{\rho(G)}\leq\frac{\Delta(G)}{\rho(G)}|U|.$
This proves that uniformly dense graphs are $(\rho(G)/\Delta(G))$-tough.
Let now $G$ be a connected, $d$-regular, uniformly dense graph. Let $U\subseteq V$ be a subset of vertices whose removal from $G$ increases the number of components, and let $T$ be a spanning tree of $G$. We write $G^{\prime}$ for the graph $G$ with vertices $U$ removed and $F^{\prime}:=\{e\in T:e\subseteq V\backslash U\}$. We note that $F^{\prime}$ is not necessarily a spanning tree or forest of $G^{\prime}$, and in general we have $c(F^{\prime})\geq c(G^{\prime})$. See Figure 6 for an example of the construction.
Figure 6: A graph $G$ and spanning tree $T\subset E(G)$. The graph $G^{\prime}$ and $F^{\prime}\subset E(G^{\prime})$ are obtained by removing the vertices in $U$ from $G$. Note that in this example $F^{\prime}$ is not a spanning forest of $G^{\prime}$, and $c(F^{\prime})>c(G^{\prime})$.
Let $\deg_{T}(v):=|\{e\in T:e\ni v\}|$ denote the degree of a vertex $v$ in $T$, and let
$\epsilon_{U}(T):=|\{e\in T:e\subseteq U\}|$
be the number of edges in $T$ between pairs of vertices in $U$. For the edge set $F^{\prime}$, we know that $|F^{\prime}|=|V\setminus U|-c(F^{\prime})$ since the set contains no cycles, and therefore we can write
00 0 $=|F^{\prime}|-|V\setminus U|+c(F^{\prime})$
$=\left[|T|-\left(\sum_{v\in U}\deg_{T}(v)-\epsilon_{U}(T)\right)\right]-\Big{[}|V|-|U|\Big{]}+c(F^{\prime})$
$=(|T|-|V|+1)-1-\sum_{v\in U}\deg_{T}(v)+\epsilon_{U}(T)+|U|+c(F^{\prime}).$
Since $G$ is connected, we have $|T|=\operatorname{rank}(G)=|V|-1$ and the equality above further simplifies to
$\sum_{v\in U}\deg_{T}(v)=\epsilon_{U}(T)+|U|+c(F^{\prime})-1.$ (5)
By uniform density, there exists a normalized $E$-uniform measure $\mu$ on the spanning trees $\mathcal{B}$ of $G$. By Lemma 2.1 this measure satisfies $\mu(\{T\in\mathcal{B}:T\ni e\})=\rho(G)^{-1}$ for every edge $e$. For any vertex $v\in V$ and summing over spanning trees, we find
$\sum_{T\in\mathcal{B}}\mu(T)(2-\deg_{T}(v))$ $=\sum_{T\in\mathcal{B}}\mu(T)\left(2-\sum_{e\in T:e\ni v}1\right)$
$=2-\sum_{T\in\mathcal{B}}\sum_{e\in T:e\ni v}\mu(T)\qquad\qquad(\mu\text{~{}is normalized})$
$=2-\sum_{e\in E:e\ni v}\sum_{T\in\mathcal{B}:T\ni e}\mu(T)\qquad(\text{switch summation order})$
[…]
$=2-\frac{d(|V|-1)}{|E|}=2-\frac{2(|V|-1)}{|V|}>0.$
Summing this inequality over the vertices in $U$, we obtain
00 0 $<\sum_{v\in U}\sum_{T\in\mathcal{B}}\mu(T)(2-\deg_{T}(v))$
$=\sum_{T\in\mathcal{B}}\mu(T)\left(2|U|-\sum_{v\in U}\deg_{T}(v)\right)$
[…]
$\leq\sum_{T\in\mathcal{B}}\mu(T)\left(|U|-c(G^{\prime})+1\right)\quad\text{~{}($\epsilon_{U}(T)\geq 0$ and $c(F^{\prime})\geq c(G^{\prime}))$}$
$=|U|-c(G^{\prime})+1.$
Since $c(G^{\prime})$ and $|U|$ are integers, the strict inequality implies $c(G^{\prime})\leq|U|$. We recall that $G^{\prime}$ is the graph $G$ with vertices $U$ removed. This completes the proof that $G$ is $1$-tough. ∎
Our proof of the $1$-tough property in Theorem 1.1 is heavily based on the proof of Theorem 3.4.18 in Fiedler’s book [11], which works with a geometric condition on graphs to guarantee the existence of a measure $\mu$ such that $\sum_{T}\mu(T)(2-\deg_{T}(v))$ is positive for each vertex $v$. This condition is satisfied for connected, regular, uniformly dense graphs.
To conclude, we highlight one implication of the connectivity Theorem 1.1 for uniformly dense graphs. A perfect matching in a graph is a subset $A\subseteq E$ of edges such that $A$ covers every vertex exactly once.
Corollary 3.13.
Let $G$ be a connected, regular, uniformly dense graph. Then, either $G$ has a perfect matching, or $G\backslash v$ has a perfect matching, for every vertex $v\in V(G)$.
Proof.  An even $1$-tough graph has a perfect matching (see Theorem 91 in [2]) and an odd $1$-tough graph has a perfect matching in every subgraph with one vertex removed (see Theorem 92 in [2]). The claim thus follows immediately from Theorem 1.1. ∎
3.4 Spectrum
In this section, we study spectral properties of uniformly dense graphs; that is, we define linear operators associated with graphs and study how uniform density is reflected in the eigenvalues of these operators. For more background on spectral graph theory, we refer to [21, 6, 4, 20, 15].
We again work with a graph $G=(V,E)$ and note that removing or adding isolated vertices to a graph does not change the density or uniform density. Thus, for a given set $A\subseteq E$ of edges, we define $G|A$ as the graph obtained from $(V,A)$ by removing all of its isolated vertices and write $V|_{A}=V(G|A)$ for the vertices of $G|A$.
Let $C(V)$ denote the vector space of functions $f:V\rightarrow\mathbb{R}$, and let $C(E)$ be the vector space of functions $\gamma:E\rightarrow\mathbb{R}$. The normalized Laplacian $L=L(G)$ associated with $G$ is the linear map $L:C(V)\rightarrow C(V)$ defined by
$Lf(v):=f(v)-\frac{1}{\deg(v)}\sum_{u:\{u,v\}\in E}f(u),$
for $f\in C(V)$ and $v\in V$. In matrix notation, this operator is usually written as $L=I-AD^{-1}$ where $I$ is the identity matrix, $A$ the adjacency matrix and $D=\operatorname{diag}(\deg(v_{1}),\dots,\deg(v_{n}))$ the diagonal degree matrix.
Now, for each edge, we fix an arbitrary orientation, that is, we let one of its endpoints be its input, and we let the other endpoint be its output. Fixing an orientation is needed in order to give the following definition, but changing the orientation of any edge does not affect the following operator.
The edge Laplacian $L^{1}=L^{1}(G)$ associated with $G$ is the linear map $L^{1}:C(E)\rightarrow C(E)$ defined by
$L^{1}\gamma(e):=\frac{\sum_{e_{1}:v\text{ input}}\gamma(e_{1})-\sum_{e_{2}:v\text{ output}}\gamma(e_{2})}{\deg v}-\frac{\sum_{e^{\prime}_{1}:w\text{ input}}\gamma(e^{\prime}_{1})-\sum_{e^{\prime}_{2}:w\text{ output}}\gamma(e^{\prime}_{2})}{\deg w},$
for $\gamma\in C(E)$ and $e\in E$ that has input $v$ and output $w$.
As shown, for instance, in [21], $L$ has $|V|$ real, nonnegative eigenvalues whose algebraic and geometric multiplicity coincide, and similarly $L^{1}$ has $|E|$ real, nonnegative eigenvalues. The nonzero eigenvalues of the two operators coincide, and they are known to encode several geometric properties of the graph associated with them. Moreover, according to the Courant–Fischer–Weyl min-max Principle, the eigenvalues of $L$ are given by minimax values of the Rayleigh quotients
$\operatorname{RQ}(f):=\frac{\sum_{\{v,w\}\in E}\biggl{(}f(v)-f(w)\biggr{)}^{2}}{\sum_{v\in V}\deg v\cdot f(v)^{2}},$
for $f\in C(V)$ [21].
A first spectral result on uniformly dense graphs follows from the fact that the multiplicity of zero as an eigenvalue of $L(G)$ is equal to the number of connected components of $G$ and the multiplicity of zero as an eigenvalue of $L^{1}(G)$ is equal to the number of cycles in $G$. Let $n_{0}(\cdot)$ denote the multiplicity of zero as an eigenvalue of a matrix, then we have [21]
$n_{0}(L(G))=c(G)\quad\text{~{}and~{}}\quad n_{0}(L^{1}(G))=\beta(G).$ (6)
The uniform density condition of a graph can be written in terms of the Laplacian spectra as follows:
Theorem 3.14.
The following are equivalent for a graph $G$:
  1. (1)
    $G$ is uniformly dense.
  2. (2)
    $\operatorname{rank}(L(G))-\operatorname{rank}(L(G|A))\leq\frac{|E|-|A|}{\rho(G)}$ for all $\emptyset\neq A\subseteq E$.
  3. (3)
    $\frac{n_{0}(L^{1}(G|A))}{|A|}\leq\frac{n_{0}(L^{1}(G))}{|E|}\text{~{}for all $\emptyset\neq A\subseteq E$.}$
Proof.  $(1)\Leftrightarrow(2).$ Note that $\operatorname{rank}(L(G))=|V|-n_{0}(L(G))$. Following (6), the difference of the Laplacian ranks can be written as
$\operatorname{rank}(L(G))-\operatorname{rank}(L(G|A))$ $=(|V|-c(G))-(|V|_{A}|-c(G|A))$
$=\operatorname{rank}(E)-\operatorname{rank}(A)$
$=c(A)-c(E).$
The rank in the first line is the linear-algebraic rank of the Laplacian matrices, while the rank in the second line is the rank in the cycle matroid $M(G)$. By Proposition 3.12 we know that $c(A)-c(E)\leq(|E|-|A|)/\rho(G)$ if and only if $G$ is uniformly dense; this proves the equivalence.
$(1)\Leftrightarrow(3).$ We recall that $\beta(G)=|E|-\operatorname{rank}(G)$. Invoking expression (6), we then find
$\frac{n_{0}(L^{1}(G|A))}{|A|}\leq\frac{n_{0}(L^{1}(G))}{|E|}$ $\iff\frac{|A|-\operatorname{rank}(A)}{|A|}\leq\frac{|E|-\operatorname{rank}(E)}{|E|}$
$\iff\frac{\operatorname{rank}(A)}{|A|}\geq\frac{\operatorname{rank}(E)}{|E|}$
$\iff\rho(A)\leq\rho(E),$
for all $\emptyset\neq A\subseteq E(G)$. By Theorem 1.1 the latter inequality is equivalent to uniform density.∎
Now, we let $\lambda_{\max}(G)$ denote the largest eigenvalue of $L(G)$, and similarly, given $A\subseteq E$, we let $\lambda_{\max}(G|A)$ denote the largest eigenvalue of $L(G|A)$. We also define the boundary of $A$ as
$\delta(A):=\{e\in E\setminus A\,:\,\text{at least one endpoint of }e\text{ is in }V|_{A}\,\}.$
Given a set $S$ of vertices, we let
$\operatorname{vol}(S):=\sum_{v\in S}\deg v$
be the volume of $S$. Moreover, given a vertex $v$, we let $\deg_{A}v$ denote the degree of $v$ in the graph $G|A$, and given $S\subseteq V|_{A}$, we let
$\operatorname{vol}_{A}(S):=\sum_{v\in S}\deg_{A}v.$
Theorem 3.15.
Let $G$ be a graph. Then
$\lambda_{\max}(G)\geq 2\cdot\rho(G)^{-1}$ (7)
and
$\lambda_{\max}(G)\geq 2\cdot\rho(A)^{-1}\frac{|A|}{|A|+|\delta(A)|}\quad\text{~{}for all $\emptyset\neq A\subseteq E$.}$ (8)
Moreover, if $G$ is uniformly dense, then
$\lambda_{\max}(G|A)\geq 2\cdot\rho(G)^{-1}\quad\text{for all }\emptyset\neq A\subseteq E.$ (9)
Proof.  Let $G$ be a graph with $n$ vertices and $m$ edges and let $T$ be a spanning tree of $G$. Then $T$ has $c(G)$ components and $n-c(G)$ edges. Furthermore, $T$ is bipartite and there exists a partition $V=V_{1}\sqcup V_{2}$ of the vertex set such that all edges of $T$ have one endpoint in $V_{1}$ and one endpoint in $V_{2}$. Hence, the number of edges between $V_{1}$ and $V_{2}$ in $G$, denoted $|E(V_{1},V_{2})|$, satisfies
$|E(V_{1},V_{2})|\geq n-c(G).$
Now, let $f:V\rightarrow\mathbb{R}$ be defined by
$f(v):=\begin{cases}1,&\text{if }v\in V_{1}\\ -1,&\text{if }v\in V_{2}.\end{cases}$
Then, by the Courant–Fischer–Weyl min-max Principle,
$\lambda_{\max}(G)\geq\textrm{RQ}(f)=\frac{4\cdot|E(V_{1},V_{2})|}{2m}\geq 2\cdot\frac{n-c(G)}{m}=2\cdot\rho(G)^{-1}.$
This proves (7). In the same way, one can show that
$\lambda_{\max}(G|A)\geq 2\cdot\rho(G|A)^{-1}=2\cdot\rho(A)^{-1},\quad\text{forall }\emptyset\neq A\subseteq E.$
Hence, if $G$ is uniformly dense, we can infer that
$\lambda_{\max}(G|A)\geq 2\cdot\rho(G)^{-1},\quad\text{for all }\emptyset\neq A\subseteq E.$
This proves (9).
Similarly, let $\tilde{T}$ be a spanning tree of $G|A$ for some $\emptyset\neq A\subseteq E$ and write $n_{A}=|V|_{A}|$ and $m_{A}=|A|$. Then $\tilde{T}$ has $n_{A}-c(G|A)$ edges and, since it is bipartite, there exists a partition $V|_{A}=V_{1}\sqcup V_{2}$ of its vertex set such that all edges of $\tilde{T}$ have one endpoint in $V_{1}$ and one endpoint in $V_{2}$. Hence, the number of edges between $V_{1}$ and $V_{2}$ in $G$ satisfies
$|E(V_{1},V_{2})|\geq n_{A}-c(G|A).$
Also,
$\operatorname{vol}(V_{1})+\operatorname{vol}(V_{2})=\operatorname{vol}(V|_{A})\leq\operatorname{vol}_{A}(V|_{A})+2|\delta(A)|=2m_{A}+2|\delta(A)|.$
Let now $f:V\rightarrow\mathbb{R}$ be defined by
$f(v):=\begin{cases}1,&\text{if }v\in V_{1},\\ -1,&\text{if }v\in V_{2},\\ 0,&\text{otherwise.}\end{cases}$
Then, as in the proof of [3, Theorem 4.5], by the Courant–Fischer–Weyl min-max Principle,
$\lambda_{\max}(G)\geq\textrm{RQ}(f)\geq\frac{4\cdot|E(V_{1},V_{2})|}{\operatorname{vol}(V_{1})+\operatorname{vol}(V_{2})}\geq\frac{2\cdot(n_{A}-c(G|A))}{m_{A}+|\delta(A)|}=2\cdot\rho(A)^{-1}\frac{m_{A}}{m_{A}+|\delta(A)|}.$
This proves (8). ∎
Example 3.16.
Let $G$ be a tree graph. In this case $\rho(G)=1$. Also, $G$ is bipartite and the same holds for $G|A$, for all $A\subseteq E$. Hence,
$\lambda_{\max}(G)=\lambda_{\max}(G|A)=2,$
and both (7) and (9) are equalities in this case.
The next example gives a graph that does not satisfy (9) and therefore is not uniformly dense.
Example 3.17.
Let $G$ be the graph given by the complete graph on $5$ vertices that shares a common edge with the cycle graph on $20$ vertices (see Figure 7). Then, $c(G)=1$, $|V|=23$, and $|E|=29$. If $A$ is the set of edges of the complete graph on $5$ vertices, then
$\lambda_{\max}(G|A)=\frac{5}{4}<2\cdot\frac{|V|-c(G)}{|E|}=\frac{44}{29}.$
This shows that $G$ does not satisfy (9). Hence, it is not uniformly dense.
Figure 7: The graph in Example 3.17.
Remark 3.18.
If $G$ is uniformly dense and $A\subseteq E$ is such that $G|A$ is obtained from $G$ by simply deleting vertices and the edges in which these vertices are contained, then by Theorem 3.15 and by the Cauchy interlacing Theorem,
$\lambda_{\max}(G)\geq\lambda_{\max}(G|A)\geq 2\cdot\rho(G)^{-1}.$
However, for a general set $A\subseteq E$, there are no interlacing results for $\lambda_{\max}(G)$ and $\lambda_{\max}(G|A)$.
Remark 3.19.
The last statement in Theorem 3.15 tells us that, the bigger the quantity $\rho(G)^{-1}$ is, and therefore the sparser $G$ is, then the bigger $\lambda_{\max}(G|A)$ must be, for all $A\subseteq E$. But, from classical results in spectral graph theory [6], the bigger $\lambda_{\max}(G|A)$ is, then the sparser $G|A$ must be, since the largest eigenvalue of the normalized Laplacian achieves its smallest possible value for complete graphs. This is in line with the interpretation of uniform density in Remark 3.1.
For uniformly dense graphs, we can strengthen the lower bound (7) on the largest Laplacian eigenvalue by making use of the following tree packing result of Catlin et al.
Lemma 3.20 ([5, Thm. 6]).
A uniformly dense graph contains $\bigl{\lfloor}\rho(G)\bigr{\rfloor}$ edge-disjoint spanning trees.
This leads to the following inequality:
Theorem 3.21.
Let $G$ be a uniformly dense graph. Then
$\lambda_{\max}(G)\geq 2\cdot\rho(G)^{-1}+2\cdot\frac{\left(\left\lfloor\rho(G)\right\rfloor-1\right)c(G)}{|E|}.$ (10)
Proof.  Let $G$ be a uniformly dense graph with $n$ vertices and $m$ edges. By Lemma 3.20, there exists a set of $\bigl{\lfloor}\rho(G)\bigr{\rfloor}\geq 1$ edge-disjoint spanning trees. Let $T$ be any of these trees, with $V=V_{1}\sqcup V_{2}$ a vertex partition such that all edges of $T$ are between $V_{1}$ and $V_{2}$ and define $f:V\rightarrow\{-1,+1\}$ as in the proof of Theorem 3.15. Then, by the Courant–Fischer–Weyl min-max Principle,
$\lambda_{\max}(G)\geq\operatorname{RQ}(f)=\frac{4\cdot|E(V_{1},V_{2})|}{2m},$
where $|E(V_{1},V_{2})|$ is the number of edges between $V_{1}$ and $V_{2}$ in $G$. Since $T$ has all $n-c(G)$ edges between $V_{1}$ and $V_{2}$ and every other spanning tree has at least $c(G)$ edges between $V_{1}$ and $V_{2}$, we have $|E(V_{1},V_{2})|\geq n-c(G)+(\lfloor\rho(G)\rfloor-1)c(G)$. This completes the proof. ∎
In general, (10) is tighter than (7) since $\lfloor\rho(G)\rfloor\geq 1$.
4 Real representable matroids
We now consider real representable matroids, which are matroids whose structure is determined by a real matrix. We show that this additional algebraic structure guarantees that certain $E$-uniform basis measures with nice properties always exist for uniformly dense matroids (Theorem 4.6). As an important practical result, we show that strict uniform density can be recognized in polynomial time for real representable matroids (Theorem 1.1). Second, we show that strictly uniformly dense real matroids can be represented by orthogonal projection matrices with constant diagonal (Theorem 1.1) and that they are parametrized by a subvariety of the Grassmannian.
4.1 Definition and examples
A matroid $M$ of size $n$ and rank $k$ is real representable if there exists a real, full-rank $k\times n$ matrix $X$ with columns indexed by $E(M)$ such that a $k$-sized subset $B\subseteq E(M)$ is a basis if and only if $\det(X_{B})\neq 0$, where $X_{B}$ is the submatrix with columns in $B$. We call $X=X(M)$ a representation of $M$ and note that representations are not unique; in particular, scaling each column of $X$ by some nonzero number gives another representation. By the construction above, every full-rank $k\times n$ matrix $X$ determines a matroid $M(X)$. The linear-algebraic rank of a matrix $X$ and a subset of columns $X_{A}$ is equal to the corresponding rank in the matroid $M(X)$.
We give a number of examples of real representable matroids.
Example 4.1.
Let $M$ be the matroid with ground set $[5]$ and bases $\{123,134,234,135,345\}$. This matroid is real representable and a matrix and vector representation of $M$ are shown in Figure 8.
Figure 8: A matrix and vector representation of the matroid in Example 4.1.
We note that scaling any of the columns of $X$ or, equivalently, any of the vectors by some positive real number $w$ still gives a valid representation. Since $3$ is a coloop and the other elements are not, $M$ is not uniformly dense.
Example 4.2 (Diagonal matrix).
A diagonal matrix $D$ is a matrix with real numbers $d_{1},\dots,d_{n}$ on the diagonal and zeroes otherwise. The rank of a diagonal matrix is equal to the number of nonzero $d_{i}$’s and it is uniformly dense if and only if all numbers are nonzero, in which case $M(D)=U_{n}^{n}$, or zero, in which case $M(D)=U_{n}^{0}$.
Example 4.3 (Incidence matrix).
Let $G=(V,E)$ be a graph and fix an arbitrary bijection $o_{e}:e=\{u,v\}\rightarrow\{-1,1\}$ for each $e\in E$. The incidence matrix of $G$ is the $|V(G)|\times|E(G)|$ matrix $\tilde{X}(G)$ with entries
$(\tilde{X}(G))_{ve}:=\begin{cases}o_{e}(v)\text{~{}if $v\in e$}\\ 0\text{~{}otherwise}.\end{cases}$
The sum over all rows corresponding to the vertices of a connected component of $G$ is zero, which means that the incidence matrix is rank deficient by $c(G)$. The reduced incidence matrix $X(G)$ is obtained by selecting one vertex in every connected component and deleting the corresponding row from the incidence matrix. This is a full-rank matrix that represents the graphic matroid of $G$, i.e. $M(X(G))=M(G)$ independent of the choice of $o_{e}$ and deleted vertices (i.e. rows); see for instance [18, 24]. Figure 9 below shows the reduced incidence matrix from the graph in Example 3.2.
Figure 9: A graph $G$ and its reduced incidence matrix $X(G)$. The row corresponding to the leftmost vertex in $G$ is removed from the incidence matrix.
We consider a second, less standard representation of matroids. A real symmetric matrix $T$ is an orthogonal projection matrix if $T^{2}=T$. An orthogonal projection matrix $T=T(M)$ is a projection representation of matroid $M$ if a $k$-sized subset $B\subseteq E(M)$ is a basis in $M$ if and only if the corresponding $k\times k$ principal submatrix $T_{BB}$ is non-singular. Projection representations are related to classical matrix representations by the identities $T=X^{T}(XX^{T})^{-1}X$ and $\operatorname{row}(X)=\operatorname{im}(T)$, where $\operatorname{row}(~{})$ denotes the row span. We also write $T=T(X)$ and $X=X(T)$ to make the relation between representations explicit.
Example 4.4.
The projection representation of the matroid in Example 3.2 and 4.3 equals
$T=\frac{1}{8}\begin{pmatrix}5&-1&-1&3&-2\\ -1&5&-3&1&2\\ -1&-3&5&1&2\\ 3&1&1&5&2\\ -2&2&2&2&4\end{pmatrix}.$
All $3\times 3$ principal minors of $T$ are nonzero, except $\det(T_{145,145})=\det(T_{235,235})=0$. These are precisely the non-bases of the matroid $M(T)$. For graphic matroids, this projection matrix is called the transfer current matrix [18], and it is a projection onto the cut-space of the graph.
4.2 Strictly uniformly dense real representable matroids
In this section we show that strictly uniformly dense real representable matroids always have an $E$-uniform basis measure of a special form (see Theorem 4.6) and a projection representation with constant diagonal. From here on, we will abbreviate (strictly) uniformly dense real representable matroid by (strictly) uniformly dense real matroid.
We will make use of the following result; see for instance [19, Cor. 8.25] and [26, Cor. 7.3.9].
Lemma 4.5.
Let the matrix $A\in\mathbb{Z}^{n\times m}$ have rank $n$ and $\mathbbm{1}\in\operatorname{row}(A)$, and let $p\in\mathbb{R}^{m}_{>0}$. If the linear system $Ay=b$ has a positive solution $y\in\mathbb{R}^{m}_{>0}$, then it has a unique positive solution $\hat{y}$ of the form
$\hat{y}_{k}=p_{k}\prod_{e=1}^{n}w_{e}^{a_{ek}},$
for $k=1,\dots,m$, for some $w\in\mathbb{R}^{n}_{>0}$.
The point $\hat{y}$ is also called the Birch point, and the prescribed form means that $\hat{y}$ lies on the positive real part of the toric variety parametrized by the integer matrix $A$. Lemma 4.5 is important in the context of maximum likelihood estimation for log-linear models in algebraic statistics [26, Ch. 7.2], and in the context of the so-called moment map, which maps torus orbits in the Grassmannian to polytopes in Euclidean space, see [14] and [19, Ch. 8.2]. Here, we use Lemma 4.5 to prove the following characterization of strictly uniformly dense real matroids in terms of special kinds of measures. Recall that a matroid is strictly uniformly dense if $(\rho^{-1},\dots,\rho^{-1})$ is an interior point of the base polytope.
Theorem 4.6.
Let $M$ be a real representable matroid. Then $M$ is strictly uniformly dense if and only if it has a representation $X$ such that $\mu_{X}:\mathcal{B}(M)\ni B\mapsto\det(X_{B})^{2}$ is an $E$-uniform basis measure.
Proof.  The converse direction holds by Theorem 1.1. We prove the forward direction. Let $M$ be a strictly uniformly dense real matroid of size $n$ and rank $k$, with $m:=|\mathcal{B}(M)|$ bases and with representation $X$. Define the $n\times m$ matrix $A$ with columns given by the indicator vectors $e_{B}$ of bases of $M$. This matrix has full rank and $\mathbbm{1}\in\operatorname{row}(A)$ since every column sums to $k=\operatorname{rank}(M)$. Define the $m\times 1$ vector $p$ with entries $p_{B}:=\det(X_{B})^{2}>0$ for each $B\in\mathcal{B}(M)$. Since $M$ is strictly uniformly dense, there exists a normalized $E$-uniform measure $\mu$ on $\mathcal{B}$ with full support. Define the $m\times 1$ vector $y$ with entries $y_{B}:=\mu(B)>0$ for each $B\in\mathcal{B}(M)$. By definition of $E$-uniform measures and Lemma 2.1, we have $Ay=\rho^{-1}\cdot\mathbbm{1}$.
Since $y$ is positive, Lemma 4.5 implies that there exists a unique positive solution to equation $A\hat{y}=\rho^{-1}\cdot\mathbbm{1}$ of the form
$\hat{y}_{B}=p_{B}\prod_{e=1}^{n}w_{e}^{a_{eB}}=\det(X_{B})^{2}\prod_{e\in B}w_{e}$
for some $w\in\mathbb{R}^{n}_{>0}$. Here we used that $a_{eB}=1$ if $e\in B$ and zero otherwise. Let $\tilde{X}$ be the matrix $X$ where the $e^{\text{th}}$ column is scaled by $\sqrt{w_{e}}$; this is still a representation of $M$ and it satisfies
$\det(\tilde{X}_{B})^{2}=\det(X_{B})^{2}\prod_{e\in B}w_{e}.$
We can thus write $\hat{y}_{B}=\det(\tilde{X}_{B})^{2}$. Now let $\mu_{\tilde{X}}(B):=\det(\tilde{X}_{B})^{2}$, then by $A\hat{y}=\rho^{-1}\cdot\mathbbm{1}$ we know that $\mu_{\tilde{X}}$ is an $E$-uniform basis measure corresponding to the representation $\tilde{X}$. This completes the proof. ∎
We will call the measure in Theorem 4.6 a determinantal measure of $M$ and write $\mu_{X}$, where $X$ is the distinguished representation guaranteed by the theorem. This terminology refers to the fact that $\mu_{X}$ is called a projection determinantal measure in the context of determinantal point processes, see for instance [18].
[…]
Proof.  This result follows directly from the theory of operator scaling. Theorem 1.2 in [12] says that there is a polynomial-time algorithm to test if a given point lies inside a so-called Brascamp–Lieb polytope, which is a polytope determined by some matrix $B$; Section 8 in [12] establishes that matroid polytopes of real matroids with representation $X$ are examples of Brascamp–Lieb polytopes, where $X$ takes the role of $B$ (this is originally a result due to Barthe [1]). ∎
We refer to [12] for further details on the precise conditions for ‘polynomial time’; in short, this means polynomial in the bit size of the representation matrix $X$. Since graphic matroids are representable with a distinguished and easy to construct representation (see Example 4.3), Theorem 1.1 implies that uniformly dense graphs can be recognized efficiently:
Corollary 4.7.
There is a polynomial-time algorithm that tests if a graph $G$ is uniformly dense and, if it is not, outputs a set of edges whose density is larger than the density of the graph.
The second lemma we use is a linear-algebraic identity that follows from the theory of determinantal point processes; see for instance [18].
Lemma 4.8.
Let $T$ be an orthogonal projection matrix of size $n$ and rank $k$. Then
$\frac{\sum_{S\in{[n]\choose k}}\det(T_{SS})e_{S}}{\sum_{S\in{[n]\choose k}}\det(T_{SS})}=\operatorname{diag}(T).$
Combining Theorem 4.6 and Lemma 4.8, we find:
[…]
Proof.  (Proof of forward direction.) Let $M$ be a strictly uniformly dense real matroid of size $n$ and rank $k$. By Theorem 4.6, $M$ has a representation $X$ such that the determinantal measure $\mu_{X}:\mathcal{B}\ni B\mapsto\det(X_{B})^{2}$ is $E$-uniform. Let $T=T(X)=X^{T}(XX^{T})^{-1}X$ be the corresponding projection representation and observe that
$\det(T_{SS})=\frac{\det(X_{S})^{2}}{\det(XX^{T})},\quad\text{for all }S\in{[n]\choose k}.$
We can thus write the normalized determinantal basis measure as $\tilde{\mu}_{X}:\mathcal{B}\ni B\mapsto\tfrac{1}{Z}\det(T_{BB}).$ By Lemma 4.8 we then have
$\operatorname{diag}(T)=\frac{\sum_{S\in{[n]\choose k}}\det(T_{SS})e_{S}}{\sum_{S\in{[n]\choose k}}\det(T_{SS})}=\sum_{B\in\mathcal{B}(M)}\tfrac{1}{Z}\det(X_{B})^{2}e_{B},$
where $Z=\sum_{B\in\mathcal{B}(M)}\det(X_{B})^{2}$. Looking at the $e^{\text{th}}$ diagonal entry, we find
$T_{ee}=\sum_{B\ni e}\tfrac{1}{Z}\det(X_{B})^{2}=\rho(M)^{-1},$
where the second equality follows from Lemma 2.1. Thus $T$ has constant diagonal equal to $\rho(M)^{-1}$ which completes the proof of the forward direction.
(Proof of converse direction.) Let $T$ be an orthogonal projection matrix with constant diagonal $k/n$ and let $\tilde{X}=\tilde{X}(T)$. Then $T$ is a projection representation of matroid $M(\tilde{X})$ with density $\rho(M(\tilde{X}))=n/k$. By the same steps as in the forward direction, the normalized measure
$\mu_{T}(B):=\tfrac{1}{Z}\det(T_{BB})=\tfrac{1}{Z}\det(\tilde{X}_{B})^{2}=\mu_{\tilde{X}}(B)$
is an $E$-uniform measure on $\mathcal{B}(M(\tilde{X}))$ and thus $M(\tilde{X})$ is strictly uniformly dense. ∎
Example 4.9.
Let $M$ be the matroid from Example 3.2. This is the rank $3$ matroid on $[5]$ with non-bases $125$ and $345$. This matroid is graphic, real representable and strictly uniformly dense. Starting from the representation in Figure 9, i.e. the reduced incidence matrix of the associated graph, we use the operator scaling algorithm to calculate the column scaling vector $w=(2,2,2,2,3)$ appearing in Lemma 4.5 and the proof of Theorem 4.6. This gives the representation
$\tilde{X}=\left(\begin{smallmatrix}-\sqrt{2}&\sqrt{2}&0&0&\sqrt{3}\\ 0&-\sqrt{2}&\sqrt{2}&0&0\\ 0&0&-\sqrt{2}&-\sqrt{2}&-\sqrt{3}\end{smallmatrix}\right)$
whose corresponding determinantal measure has $\mu_{\tilde{X}}(B)=12$ for all bases that contain element $5$ in the ground set (the diagonal edge in the corresponding graph) and $\mu_{\tilde{X}}(B)=8$ otherwise. This is an $E$-uniform basis measure for $M$. The projection matrix onto $\operatorname{row}(\tilde{X})$ equals
$T(\tilde{X})=\frac{1}{10}\begin{pmatrix}6&-1&-1&4&-\sqrt{6}\\ -1&6&-4&1&\sqrt{6}\\ -1&-4&6&1&\sqrt{6}\\ 4&1&1&6&\sqrt{6}\\ -\sqrt{6}&\sqrt{6}&\sqrt{6}&\sqrt{6}&6\end{pmatrix}$
This matrix has constant diagonal $3/5$ and its non-singular $3\times 3$ principal submatrices correspond to the bases of $M$. This is the projection representation from Theorem 1.1.
Finally, we note an application of Theorem 1.1 to the properties of projection matrices.
Corollary 4.10.
Let $T$ be a constant-diagonal orthogonal projection matrix of size $n$ and rank $k$. Then for all $0<m<n$, every principal $m\times m$ submatrix $T^{\prime}$ of $T$ has $\operatorname{rank}(T^{\prime})\geq m\cdot k/n$.
Proof.  Let $T$ be a constant-diagonal orthogonal projection matrix of size $n$ and rank $k$. Following Theorem 1.1, let $M=M(T)$ be the strictly uniformly dense real matroid represented by $T$ and $X=X(T)$ the corresponding standard representation. Note that $T_{AA}=X_{A}^{T}(XX^{T})^{-1}X_{A}$ for all $A\subseteq[n]$, which implies
$\operatorname{rank}(T_{AA})=\operatorname{rank}(X_{A})=\operatorname{rank}_{M}(A)\geq\operatorname{rank}_{M}(E)\frac{|A|}{|E|}=k\cdot|A|/n\text{~{}for all $\emptyset\neq A\subset E(M)$}.$
Here, the first two ranks are linear-algebraic, while the third and fourth rankM are the rank function in the matroid $M$. The inequality follows from uniform density of $M$. ∎
4.3 The variety of uniformly dense real matroids
As discussed earlier, a full-rank $k\times n$ matrix $X$ determines a real representable matroid $M(X)$. The set of all such matrices modulo elementary row operations is called the real Grassmannian $\operatorname{Gr}(n,k)$; equivalently, this is the set of $k$-dimensional subspaces of $\mathbb{R}^{n}$. Points in the Grassmannian parametrize the real representable matroids. The embedding $p:\operatorname{Gr}(n,k)\hookrightarrow\mathbb{P}^{{n\choose k}-1}$ maps a matrix $X\in\operatorname{Gr}(n,k)$ onto its $k$-minors $p_{B}(X):=\det(X_{B})$ for each $B\in{[n]\choose k}$, up to a common non-zero scaling of all minors. These minors $p_{B}(X)$ are called the Plücker coordinates of $X$, and the vanishing of certain quadratic polynomials in these coordinates (the Plücker relations) defines the Grassmannian as a variety in ($\mathbb{P}^{{n\choose k}-1}$; see for instance [19, Ch. 5] for more details.
Following Theorem 4.6, every strictly uniformly dense real matroid $M$ has a distinguished representation $X$ whose Plücker coordinates ($p_{B}(X)=\sqrt{\mu_{X}(B)}$) satisfy the quadratic relations
$\sum_{B\ni e}p^{2}_{B}(X)=\sum_{B\ni e^{\prime}}p^{2}_{B}(X),\quad\text{for all }e\neq e^{\prime}\in E(M).$
This means that every strictly uniformly dense real matroid has a representation in the following subvariety of the Grassmannian:
$\mathcal{V}(n,k):=\left\{p\in\operatorname{Gr}(n,k)\,:\,\sum_{e\in B\in{[n]\choose k}}p_{B}^{2}-\sum_{(e+1)\in B\in{[n]\choose k}}p_{B}^{2}=0\text{~{}for$e=1,\dots,n-1$}\right\}\subseteq\mathbb{P}^{{n\choose k}-1}.$
Conversely, every point in $\mathcal{V}(n,k)$ corresponds to a matrix $X$ whose non-zero Plücker coordinates determine the bases of a strictly uniformly dense real matroid $M(X)$. In other words, the variety $\mathcal{V}(n,k)$ parametrizes the strictly uniformly dense real matroids.
A different description follows from Theorem 1.1. Let $T$ be an $n\times n$ symmetric matrix with constant diagonal $k/n$. This is a projection matrix if and only if it satisfies $T^{2}-T=0$; in other words, it is defined by the vanishing of (${n+1\choose 2}$ quadratic polynomials in its (${n\choose 2}$ off-diagonal entries $\{t_{ij}\}_{1\leq i<j\leq n}$. Another description of $\mathcal{V}(n,k)$ is thus
$\mathcal{V}(n,k)=\left\{t\in\mathbb{R}^{{n\choose 2}}\,:\,T^{2}-T=0\right\},$
where $T$ is the symmetric matrix with constant diagonal $k/n$ and off-diagonal entries $t_{ij}$. Every point in $\mathcal{V}(n,k)$ is the projection representation of a strictly uniformly dense real matroid. The algebro-geometric properties of the variety $\mathcal{V}(n,k)$ are further studied in [8].
Acknowledgements
We are grateful to Bernd Sturmfels, Benjamin Schröter and Harry Richman for helpful comments, discussions and suggestions. We thank an anonymous referee for their feedback on an earlier version of this article and for suggestions that led to Theorem 1.1.